OpenStudy (anonymous):

3y - 2(y + 5) - 6 (dived by) 2

OpenStudy (misty1212):

HI!! what is divided by two?

OpenStudy (anonymous):

The divided by symbol wouldnt show up so I replaced it with the words.... So its just part of the equation 3y - 2(y + 5) - 6 ÷ 2

OpenStudy (misty1212):

\[\frac{3y-2(y+5)-6}{2}\]

OpenStudy (misty1212):

\[\frac{3y-2y-10-6}{2}\]

OpenStudy (misty1212):

or \[\frac{y-16}{2}\]

OpenStudy (anonymous):

Okay. But that cant be simplified and the answers are simpler.

OpenStudy (anonymous):

Why would you assume its over two @misty1212 ?

OpenStudy (anonymous):

First you would combine like terms then parenthesis... I dont get why it would be over 2. It isnt [3y-2(y=5)-6]/2